The index of refraction of a substance is 1.96. What is the speed of light, in 108 m/s, within the medium?
Answer #1

Speed of light= 3*10^8 m/s

Speed of light in the given medium= 3*10^8/1.96

V= 1.53*10^8 m/s

1.53 in 10^8